# Lumina[Uncaptioned image]: Detecting Hallucinations in RAG System with Context–Knowledge Signals

### v2

@@ Appendix C Details of Datasets
The RAGTruth dataset is a human annotated hallucination detection dataset, containing 15,090 training data and 2,700 testing data. Each data point consists of a query, retrieved documents, LLM-generated answer, and span-level hallucination annotation. The dataset covers three tasks, including summarization, data to text generation, and question answering. For each query-and-documents pair, RAGTruth provides answers generated by six different LLMs, including Llama2-7B, Llama2-13B, Llama2-70B, Mistral-7B, GPT-3.5, and GPT-4. In our experiment, we also utilize the extended test set provided by Sun et al. (2025b), who curated and annotated Llama3-8B generated responses.
@@
HalluRAG is an LLM annotated hallucination detection dataset for question answering. Ridder & Schilling (2025) prompted GPT-4o to generate question given sentences from Wikipedia, then used Llama2-7B, Llama2-13B, and Mistral-7B to generate answer for each question given the relevant Wikipedia article. The hallucination labels were assigned by GPT-4o with a Chain-of-Thought (CoT) prompt and verified by human. HalluRAG contains both answerable and unanswerable questions, while we only use the answerable instances for evaluation.
@@ Appendix D Implementation Details of Lumina
$\mathcal{I}_{p_{\theta}}(a_{t}|q,d,a_{<t}):=\frac{p_{\theta}(a_{t}|q,d,a_{<t})}{p_{\theta}(a_{t,1}|q,d,a_{<t})}\cdot\mathcal{R}_{p_{\theta}}^{1}(q,d,a_{<t}).$ (11)
@@ E.2 Compare with supervised baselines
We further compare Lumina with SAPLMA (Azaria & Mitchell, 2023), a supervised approach that trained a MLP model over the internal hidden states of the last generated token to classify whether the generation is hallucination or not. Following the original paper, we use hidden states at the 20th layer as input features of SAPLMA. Result in Table 4 shows that Lumina has a competitive performance against SAPLMA and even sometimes outperforms it. Note that Table 4 doesn’t show the result of Llama3-8B as the training set doesn’t contain responses generated by Llama3-8B.
@@ Appendix F Error Analysis
To analyze the failure of Lumina, we sample 20 cases from the RAGTruth dataset that are (1) hallucinated with high-external context and low-internal knowledge scores (i.e., false negative) or (2) non-hallucinated with low-external context and high-internal knowledge scores (i.e., false positive). We qualitatively analyze these cases and categorize them into three groups:
@@ Appendix G Computational Resources
Lumina is a lightweight and efficient approach, which requires only two forward passes to obtain the necessary information to compute external context and internal knowledge scores. As Lumina does not require generating multiple samples nor training, it is easy to scale up to a large amount of data. All the experiments of Lumina are conducted on a single Nvidia H100 GPU. The execution time of computing both external context and internal knowledge scores varies depending on the length of the response. For responses around 150 tokens, the average computational time is less than 1 second.

### v3

@@ Appendix C Details of Datasets
The RAGTruth dataset is a human annotated hallucination detection dataset, containing 15,090 training data and 2,700 testing data. Each data point consists of a query, retrieved documents, LLM-generated answer, and span-level hallucination annotation. The dataset covers three tasks, including summarization, data to text generation, and question answering. For each query-and-documents pair, RAGTruth provides answers generated by six different LLMs, including Llama2-7B, Llama2-13B, Llama2-70B, Mistral-7B, GPT-3.5, and GPT-4. In our experiment, we also utilize the extended test set provided by sun2025redeep, who curated and annotated Llama3-8B generated responses.
@@
HalluRAG is an LLM annotated hallucination detection dataset for question answering. ridder2025halluragdatasetdetectingcloseddomain prompted GPT-4o to generate question given sentences from Wikipedia, then used Llama2-7B, Llama2-13B, and Mistral-7B to generate answer for each question given the relevant Wikipedia article. The hallucination labels were assigned by GPT-4o with a Chain-of-Thought (CoT) prompt and verified by human. HalluRAG contains both answerable and unanswerable questions, while we only use the answerable instances for evaluation.
@@ Appendix D Implementation Details of Lumina
$\mathcal{I}_{p_{\theta}}(a_{t}|q,d,a_{<t}):=\frac{p_{\theta}(a_{t}|q,d,a_{<t})}{p_{\theta}(a_{t,1}|q,d,a_{<t})}\cdot\mathcal{R}_{p_{\theta}}(q,d,a_{<t}).$ (11)
@@ E.2 Compare with supervised baselines
We further compare Lumina with SAPLMA (azaria-mitchell-2023-internal), a supervised approach that trained a MLP model over the internal hidden states of the last generated token to classify whether the generation is hallucination or not. Following the original paper, we use hidden states at the 20th layer as input features of SAPLMA. Result in Table 4 shows that Lumina has a competitive performance against SAPLMA and even sometimes outperforms it. Note that Table 4 doesn’t show the result of Llama3-8B as the training set doesn’t contain responses generated by Llama3-8B.
@@ E.3 Performance with Hyperparameter Tuning
E.4 Additional Ablation study
Impact of MMD approximation.
When implementing Lumina, we approximate MMD with the top $k$ tokens and set $k=100$, aiming to balance between computational cost and approximation error. To test the impact of $k$, we ablate on $k\in\{50,100,500\}$ and evaluate on the RAGTruth dataset. The result shows a consistent AUROC across different $k$, with a $0.02\%$ difference, suggesting that Lumina is insensitive to the choice of MMD approximation. Additionally, in cases where the computational power is limited, choosing $k=50$ is also considerable.
Impact of random documents.
In Section 4.3, we study the impact of noises in the retrieved and random documents. To further examine the impact of different random documents on the performance, we select 5 different random documents and use each of them to compute the hallucination score. Experiments on the RAGTruth dataset shows that the standard deviation across the 5 rounds with different random documents is <0.0025<0.0025, suggesting that Lumina is very robust to the choice of random document.
Impact of the components of the information processing rate.
Our proposed information processing rate consist of two components: layer weighting probability ratio (numerator) and entropy normalization (denominator). Table 5 shows the AUROC of ablating these two components. The result shows that both components contribute to the overall performance, justifying our design choice.
LLM Layer weighting Entropy normalization Both
Llama2-7B 0.7023 0.7164 0.7652
Llama2-13B 0.8007 0.8433 0.8554
Llama3-8B 0.7512 0.7683 0.7697
Mistral-7B 0.6111 0.6723 0.7679
Table 5: Both components of information processing rate are important. We report the AUROC of each component, as well as the performance of their combination.
@@ Appendix F Error Analysis
To analyze the failure of Lumina, we sample 40 cases from the RAGTruth dataset that are (1) hallucinated with high-external context and low-internal knowledge scores (i.e., false negative) or (2) non-hallucinated with low-external context and high-internal knowledge scores (i.e., false positive). We qualitatively analyze these cases and categorize them into three groups:
@@
We extend the error analysis by sampling 50 false positive and 50 false negative cases, and prompting GPT-5 to classify the reason for error. The result in Table 6 shows that while there are edge cases that Lumina can not handle correctly, many of the errors are due to incorrect labels and low quality of retrieved documents. For those edge cases, we observe that they usually happen when the internal knowledge and external context scores are close and when the task is more reasoning intensive. Thus, when deploying Lumina, controlling the balance between internal knowledge score and external knowledge score according to the task might be a good practice to further increase the performance.
Error Type Proportion
False Positive
Incorrect labels 32%
Low quality of retrieved documents 24%
Others 44%
False Negative
Incorrect labels 16%
Low hallucination score for summarization task 64%
Others 20%
Table 6: The errors of LUMINA are mainly due to incorrect labels, quality of retrieved documents, and task-dependent biases. We report the proportion of each error type classified by GPT-5.
@@ Appendix G Computational Resources
Lumina is a lightweight and efficient approach, which requires only two forward passes to obtain the necessary information to compute external context and internal knowledge scores. As Lumina does not require generating multiple samples nor training, it is easy to scale up to a large amount of data. All the experiments of Lumina are conducted on a single Nvidia H100 GPU. The execution time of computing both external context and internal knowledge scores varies depending on the length of the response. For responses around 150 tokens, the average computational time is less than 1 second. In addition, while Lumina requires two forward passes to compute the score, it is consistently more efficient than ReDeEP, as shown in Table 7. We believe that it is because for the external context score, ReDeEP has to store the entire attention map for every layer and use that to select the top $k$ tokens from the external context. And for the internal knowledge score, ReDeEP has to apply the logit lens before and after FFN for each transformer layer. In contrast, our external context score only requires approximating MMD at the output layer, and our internal knowledge score needs applying the logit lens only once per layer. These design choices reduce the computational cost, making Lumina much more efficient.
LLM ReDeEP Lumina
Llama2-7B 0.86 0.69
Llama2-17B 1.17 0.88
Llama3-8B 1.13 0.58
Mistral-7B 0.72 0.54
Table 7: Lumina is more effecient than ReDeEP. We report the average computational time (second/sample) for ReDeEP and Lumina.